OpenStudy (anonymous):

How do you simplify this equation? I saw it in a workbook and I can't figure out how you can simplify it because of the subtraction operation. \[\left( a - a^{-3} \right) \div \left( a^{-3} - 1 \right)\]

OpenStudy (anonymous):

youwrite it as a fraction then you flip the negative powers

OpenStudy (anonymous):

\[(a-1/a ^{3})/(1/a ^{3} -1)\]

OpenStudy (anonymous):

sorry west thats wrong

OpenStudy (anonymous):

he didnt even do anything but invert the exponents lol

OpenStudy (anonymous):

yea i think just flip the -exponets then you are done

OpenStudy (anonymous):

Then, you can easily carry out the basic operations and you get the following: \[-a ^{3}-1\]

OpenStudy (anonymous):

\[\frac{a^4-1}{1-a^3}=\frac{(a-1)(a+1)(a^2+1)}{(1-a)(1+a+a^2)}\]

OpenStudy (anonymous):

Huh. What puzzled me was that the workbook was multiple choice and the choices were stuff like just [ a ] and [ 1/a ]. Could be that I'm recalling the actual equation incorrectly since it was some time ago, but the basic premise is the same. I think I get what Aron is saying now. Cross multiplication, right? NotSObright's answer is also a neat trick that I keep forgetting when simplifying. Thanks guys!